For the following items , find the sinking fund payment

2) Camilla wats to accumulate ₱80000 in 5 years for a trip to hongkong. If her bank is paying 12% interest compounded quarterly, how much must c camilla deposit at the end of each 3 month period to reach her desired goal?

3) what sinking fund payment is required athe end of each month , at 18%. Interest compounded monthly, to amount to ₱40000 in 6 years.

4) Caspar is planning a vacation in bangkok in 4years and will need ₱75,000 for the trip .Caspar decides to set up a sinking fund savings àçcount

For the vacation . He intends to. Make regular payments at the end of each 3month period into the account that pays 6% interest compounded quarterly.what periodic sinking fund payment will allow him to achieve his vacation goal?








U
















1
Expert's answer
2021-05-18T15:15:03-0400

2)

"R=\frac{FVr+1}{(1+r)^n}=\frac{80000\cdot0.12+1}{1.12^{4\cdot5}}=P\ 995.30"


3)

"R=\frac{FVr+1}{(1+r)^n}=\frac{40000\cdot0.18+1}{1.18^{6\cdot12}}=P\ 0.05"


4)

"R=\frac{FVr+1}{(1+r)^n}=\frac{75000\cdot0.06+1}{1.06^{4\cdot4}}=P\ 1771.80"